Correct is given, please explain why for each

Concept Question 4 – An object is initially at rest. A net force (which always points in the same direction) is applied to the object so that the power supplied by the net force is constant. As the object gains speed, a) the magnitude of the net force decreases. b) the magnitude of the net force increases. c) the magnitude of the net force remains constant. d) the magnitude of the net force first increases, then remains constant. e) Not enough information is given to decide.

A is correct

Concept Question 5 – Two blocks are connected as shown. When released, the 6.00-kg block accelerates downward and the 8.00-kg block accelerates to the right. After each block has moved 2.00 cm, the total work done on the 8.00-kg block is a) greater than the total work done on the 6.00-kg block. b) the same as the total work done on the 6.00-kg block. c) less than the total work done on the 6.00-kg block, but not zero. d) zero. e) Not enough information is given to decide.

A is correct

Concept Question 6 – Rank the following objects in order of the magnitude of the momentum of the object, from largest to smallest. i. Mass = 2.0 kg, kinetic energy = 2.0 J ii. Mass = 1.0 kg, kinetic energy = 2.0 J iii. Mass = 2.0 kg, kinetic energy = 4.0 J iv. Mass = 4.0 kg, kinetic energy = 4.0 J

B is correct
Answer #1

Concept Question 4

We know that

power = force * speed

here, power is constant , only speed is changing

so,

speed = power / force

there is an inverse relation between speed and force ( if one increases, the other will decrease and vice versa)

Therefore,  As the object gains speed, the net force decreases.

------------------------------------------------------

Concept Question 5

The 6 kg block will move down. There are two forces acting on 6 kg block - one is tension ( acting upwards) and other is gravity ( acting downwards)

so,

work done on 6 kg block = T * d * cos 180 + W *d * cos 0 = 0

therefore,

total work done on 8 kg block is greater than total work done on 6 kg block

----------------------------------------------------------------------

Concept Question 6

K.E = p2 / 2m

where p is momentum

so,

p = sqrt ( 2m * K.E)

Mass = 2.0 kg, kinetic energy = 2.0 J >>>> p = 2.828 kg.m/s

Mass = 1.0 kg, kinetic energy = 2.0 J >>>> p = 2 kgm/s

Mass = 2.0 kg, kinetic energy = 4.0 J >>>> p = 4 kg.m/s

Mass = 4.0 kg, kinetic energy = 4.0 J >>>> p = 5.656 kg.m/s

just give the ranking

iv > iii > i > ii